Circular No. 6306

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) T. Zwitter and H. Mikuž, University of Ljubljana; and U. Munari and

M. Rejkuba, Asiago Observatory, communicate: “A spectrogram (range 340–910 nm) of C/1996 B2 was obtained at the European Southern Observatory’s 1.5-m telescope (+ Boller & Chivens spectrograph + CCD) on Feb. 8.2 UT. The slit (width 2′′, length 14′′) was positioned on the comet nucleus. The reflected solar spectrum accounts for the vast majority of the recorded flux, with all major absorption lines markedly visible. The strongest cometary emission is from CN(0,0) at 388 nm, followed by the C_{3} complex at 405 nm. The strongest C_{2} Swan band is that of (0,0) at 516 nm, followed by that of (1,0) at 474 nm, with that of (2,0) at 438 nm being barely detectable. The CN(0,1) 422-nm band is in moderate emission, while no cometary NaI D emission is observed.”

Corrigendum. On IAUC 6303, T. Kojima was using a 0.25-m f/6.0 reflector.

COMET C/1996 B1 (SZCZEPANSKI)

Ephemeris from orbital elements on MPEC 1996-C02:

1996 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Feb. 7 13^{h}25.^{m}41 +51^{o}15.^{′}3 0.681 1.451 119.^{o}9 36.^{o}1 8.3
12 12 58.54 +49 04.5 0.630 1.452 126.4 33.1 8.1
17 12 28.30 +45 54.6 0.587 1.458 134.0 29.2 8.0
22 11 56.43 +41 30.8 0.553 1.466 142.3 24.4 7.9
27 11 25.19 +35 50.7 0.533 1.478 150.7 19.1 7.8
Mar. 3 10 56.64 +29 10.7 0.528 1.493 157.6 14.7 7.9
8 10 32.04 +22 04.1 0.541 1.512 159.9 13.1 8.0
13 10 11.82 +15 08.2 0.570 1.533 156.5 15.0 8.1
18 9 55.82 + 8 50.7 0.613 1.557 150.1 18.6 8.4
23 9 43.57 + 3 24.4 0.669 1.584 143.2 22.2 8.6
28 9 34.52 − 1 09.9 0.735 1.613 136.6 25.2 8.9
Apr. 2 9 28.12 − 4 57.8 0.808 1.645 130.6 27.5 9.2
7 9 23.90 − 8 07.1 0.887 1.678 125.3 29.1 9.5

Total visual magnitude estimates: Feb. 1.17 UT, 9.0 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector, coma diameter 4′.5); 1.19, 8.4 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 2.19, 9.0 (Hornoch).

1996 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green